HETAV PATEL

Grade 11,

what will be the duration of day and night in hour if the diameter of earth is reduced to half of its original value the mass remain constant?a. 12 b . 6 c. 3 d. 2

what will be the duration of day and night in hour if the diameter of earth is reduced to half of its original value the mass remain constant?a. 12 b . 6 c. 3 d. 2

Grade:12th pass

1 Answers

Arun
5 years ago
moment of inertia(I) is proportional to radius^2 ang. vel (w1) = 2pi/24 i.e pi/12 Let r1 =r r2 =0.5r By law of conservation of angular momentum, I1x w1=I2xw2 I1/I2 = w2/w1 r^2/(0.5 r)^2 = w2/(pi/12) w2=pi/3 so, time period = 6hours